OpenStudy (anonymous):

Which is the area of a circle with a radius of 2 units? (what does 2 units mean)

OpenStudy (jackellyn):

units are the measurements like mm, cm, ft, yd

OpenStudy (anonymous):

2 units could be anything.. 2 meter, 2 cm

OpenStudy (anonymous):

but like i did pi r^2 and like it gave me a answer that was not even a choice !

OpenStudy (anonymous):

What you got and what are the choices?

OpenStudy (anonymous):

A.6.28 square units b.12.56 square units c.25.12 square units d.40 square units

OpenStudy (anonymous):

and i got like 10.266

OpenStudy (anonymous):

redo..:)

OpenStudy (anonymous):

Option B is the answer

OpenStudy (jackellyn):

A=pi*r^2 A= pi * (2)^2 A= pi * 4 A= 12.56 units squared

OpenStudy (anonymous):

ohh ! okay yes that is right, i was mutiplying the wrong number !

OpenStudy (anonymous):

3.14x 4= ? 3 x4= 12